# Segment Every Out-of-Distribution Object

### v3

@@ Abstract
Semantic segmentation models, while effective for in-distribution categories, face challenges in real-world deployment due to encountering out-of-distribution (OoD) objects. Detecting these OoD objects is crucial for safety-critical applications. Existing methods rely on anomaly scores, but choosing a suitable threshold for generating masks presents difficulties and can lead to fragmentation and inaccuracy. This paper introduces a method to convert anomaly Score To segmentation Mask, called S2M, a simple and effective framework for OoD detection in semantic segmentation. Unlike assigning anomaly scores to pixels, S2M directly segments the entire OoD object. By transforming anomaly scores into prompts for a promptable segmentation model, S2M eliminates the need for threshold selection. Extensive experiments demonstrate that S2M outperforms the state-of-the-art by approximately 10% in IoU and 30% in mean F1 score, on average, across various benchmarks including Fishyscapes, Segment-Me-If-You-Can, and RoadAnomaly datasets.
@@ 1 Introduction
Semantic segmentation, a critical task in computer vision, bears substantial significance acrossStatic Dataset.png various applications such as autonomous driving and aerial imagery analysis [18, 48]. While current semantic segmentation models demonstrate impressive performance, their practical deployment remains challenging. A significant obstacle lies in their limited ability to detect out-of-distribution (OoD) objects. Specifically, these models often assign pixels within the object to one of the categories used in training, leading to inaccurate segmentation masks [32, 44, 26]. Addressing the OoD detection problem in semantic segmentation is important, given that inaccuracies in OoD objects masks can lead to erroneous conclusions, ultimately posing safety concerns in applications such as autonomous driving [48, 14, 37, 4, 3].
@@
Existing OoD detection methods in semantic segmentation address this issue by assigning an anomaly score for each pixel [24, 32, 44, 35, 2]. The pixels with high anomaly scores will be considered as part of OoD objects. The anomaly scores are typically derived by the probabilistic predictions of each pixel made by the segmentation model [22, 25, 36]. For example, PEBAL [44] proposes a pixel-wise energy-based segmentation method to compute the anomaly score. RPL [32] introduces a residual pattern learning module and computes the anomaly score by energy-based method, effectively enhancing the model’s sensitivity to OOD pixels without compromising the segmentation performance on in-distribution (ID) data.
@@
While the anomaly score-based OoD detection methods can accurately identify anomalous pixels, they lack an effective way for segmenting the entire OoD object. In particular, to derive a mask for OoD objects, it is crucial to use a carefully chosen threshold to distinguish between anomalous and normal pixels [45, 9, 16]. However, determining the optimal threshold in practical applications can be a difficult task. From the Fig. 2, it is evident that the optimal threshold range for existing OoD detection methods is quite narrow; even a slight deviation, either too high or too low, can result in inaccurate segmentation. Choosing the optimal threshold often requires a dedicated validation dataset for fine-tuning the threshold. In practice, such as validation dataset is often not available. Even with the optimal threshold, the generated masks from anomaly score-based OoD detection methods may still require refinement, as the anomaly scores for certain pixels might be inaccurate, leading to fragmented or discontinuous masks. Fig. 1 illustrates various examples of masks generated by PEBAL and RPL, two state-of-the-art anomaly score-based OoD detection methods. The fragmented masks can hardly be useful in practice as the model users cannot accurately locate the OoD objects.
@@
In this paper, we propose method to convert anomaly Score To segmentation Mask, called S2M, as a simple and general OoD detection framework in semantic segmentation. Specifically, rather than deriving masks for OoD objects through thresholding anomaly scores, S2M aims to segment the entire OoD object. Thus S2M effectively mitigates the fragmentation issue associated with employing a threshold. In more details, given any anomaly scores generated by existing OoD detection methods, S2M generates box prompts from them using a prompt generator. The generated box prompt will approximately locate the OoD objects. Then, the box prompts are used as input for a promptable segmentation model to generate the masks for the OoD objects. Fig. 1 shows several examples of masks generated by S2M for the OoD objects. Different from using a threshold to generate the masks, the masks generated by S2M accurately segment the entire OoD objects while avoiding the inclusion of normal pixels in the mask. Compared with existing OoD detection methods, S2M has several advantages: 1) Simple: S2M is a simple pipeline that is easy to train and deploy, requiring no hyperparameter tuning. 2) General: S2M can be integrated with any anomaly score-based OoD detection methods to generate high-quality OoD masks. 3) Effective: S2M can accurately segment the OoD objects without creating fragmented and inaccurate masks.
@@
  • We extensively evaluate S2M on commonly used OoD segmentation benchmarks including Fishyscapes, Segment-Me-If-You-Can, and RoadAnomaly datasets, the results show that S2M improves the state-of-the-art by approximately 10% in IoU and 30% in mean F1 score, on average, across all benchmarks.
@@ 2 Related work
Recent trends in semantic segmentation have shifted towards transformer-based architectures and attention mechanisms, which offer improved handling of contextual relationships within images [13, 50, 43]. In contrast to these typical segmentation approaches, the prompt-based segmentation presents a unique paradigm. Its primary strength lies in its robust segmentation capabilities, enabling the efficient delineation of objects within an image. For example, CLIPSeg [34] introduce a system capable of generating image segmentation at test time based on any given prompt, whether text or image, thereby enabling a unified model for three distinct segmentation tasks. Segment Anything Model (SAM) [26] is known for its robust capability to generate precise masks for every object, demonstrating exceptional performance in object segmentation tasks.
@@ 3.2 Anomaly Score to Box Prompt
Figure 4: Box prompts can lead to more accurate segmentation for the OoD objects compared to point prompts. We derive point prompts from the locations corresponding to the extreme values of the anomaly scores. Visual analysis indicates that box prompts substantially improve the model’s tolerance to noise. We employ all generated box prompts to obtain a comprehensive mask of the OoD object, which ensures that the entire object is covered, rather than just a portion of it.
@@ 3.3 Outlier Exposure
In the OE process, a transformation function $T$ is applied to the OoD object masks $y^{out}$. This function adjusts the size and position of the OoD objects masks to align with the $\mathcal{D}^{in}$ image dimensions. The OE process is mathematically formulated as follows:
@@ 3.4 Prompt Generator Training
The initial step of training the prompt generator involves transforming the mask labels $y^{oe}$ into prompt labels. In particular, a transformation function $T_{box}$ is used to generate box prompt $B_{prompt}$ from $y^{oe}$,
@@ 3.5 Inference Pipeline
During inference, the input image is processed using a state-of-the-art OoD detection method, to compute an anomaly score map. This map is then fed into the prompt generator, which yields bounding box prompts indicating the locations of potential OoD objects. Subsequently, we employ a promptable segmentation model which takes both the prompts and the original image as inputs and produces masks of the OoD objects. The inference pipeline is shown in Fig. 5.
@@ 4 Experiments
DenseHybrid [17] 22.80 41.93 31.30 10.93 25.14 15.37 30.32 45.10 41.31 23.54 40.25 31.85 19.50 27.22 29.33
Synboost [11] 22.85 41.94 31.34 11.45 25.14 15.37 23.54 40.24 31.85 25.15 10.93 15.37 19.50 27.22 29.33
PEBAL [44] 23.65 69.00 29.66 7.03 28.21 9.70 21.5 44.47 36.35 4.12 27.25 6.27 15.51 33.80 23.87
RPL+CoroCL [32] 17.43 71.73 22.78 5.52 29.83 7.60 22.95 70.98 32.44 13.01 65.82 16.64 16.83 51.00 24.57
S2M (Ours) 72.54 81.98 76.57 35.52 34.93 41.52 58.55 85.06 65.13 64.80 69.63 68.73 49.09 58.58 55.73
Table 1: S2M outperforms the state-of-the-art methods on Fishyscapes Static (FS Static), Fishyscapes Lost and Found (FS Lost&Found), SMIYC Anomaly, SMIYC Obstacle and RoadAnomaly validation set. For our method, the IoU is calculated without thershold and using all produced prompt boxes. Both AuIoU and mean F1 scores are calculated by iterating through all thresholds. To facilitate a fair comparison with existing methods, we use the confidence of the box prompt as the anomaly score when calculating AuIoU and mean F1.
@@ 4.1 Experimental Setup
  • PEBAL [44]. This method introduces pixel-wise energy-biased abstention learning, a method that synergistically optimizes a novel pixel-wise anomaly abstention learning framework along with energy-based models.
@@ 4.2 Main Results
The results also show that existing anomaly score-based OoD detection methods, such as DenseHybrid, RPL, and PEBAL, are less effective in producing accurate masks for the OoD objects. While the RPL+CoroCL method achieves high IoU values, it does not perform as well across other metrics, such as the mean F1 score [5], revealing a shortfall in achieving highly accurate masks for OoD objects [5]. Our method, however, excels in all assessed metrics. The IoU achieved by our method surpasses that of the state-of-the-art methods by approximately 10% across all datasets, on average. Specifically, on the Fishyscapes Lost and Found validation dataset, our method outperforms RPL by a significant margin of 14.08%.
@@
0% 77.23 61.74 66.02 67.02 55.15 52.13
1% 79.20 61.06 71.70 72.42 59.18 59.13
2% 85.06 65.13 69.63 68.73 58.58 55.73
3% 77.84 53.01 59.53 62.34 56.28 53.83
4% 73.89 54.44 65.83 63.34 56.60 55.86
Table 2: Performance across different noise intensities. We introduced a random perturbation to the magnitude of the anomaly scores and observed that a fluctuation of 1% yielded the best training outcomes on more datasets. All experimental results reflect the performance after 80 epochs of training. The mean F1 scores are derived by systematically iterating through all thresholds in increments of 0.01.
@@ 4.3 Ablation Studies
Generalize to other anomaly scores. To further validate the generalization of our method, we leverage PEBAL [44] for generating the anomaly scores in S2M. Applying PEBAL to generate anomaly scores for the synthetic training dataset and utilizing the previously mentioned training strategy, we also observed similar performance improvement. Table 4 shows that a significant increase in terms of IoU, with improvements of 19.83% and 21.83% on the SMIYC anomaly and obstacle tracks [5] compared with PEBAL , respectively. This shows that S2M can generalize to various anomaly scores and be easily integrated with existing anomaly score-based OoD detection methods.
Data augmentation. We experimented with different levels of random fluctuations within a specific range to the values of the anomaly scores and subsequently tested the model’s performance on the Fishyscapes Static and Fishyscapes Lost & Found [3] datasets. Table 2 indicate that the model achieves optimal overall performance with the addition of 2% noise. The model exhibited a 3.31% increase in the best IoU on the Fishyscapes Static dataset, and there was also an improvement in the mean F1 scores across both datasets. On the Lost & Found dataset, the best IoU experienced a marginal sacrifice of only 0.25%. Furthermore, it was observed that an excessive noise level leads to a significant reduction in the mean F1 score, suggesting that overly strong noise can adversely affect model training, hindering the model’s ability to accurately generate prompts.
Different choices of SAM. We investigated various SAM backbone configurations, including ViT-B, ViT-L, and ViT-H [26], the corresponding S2M is denoted as S2M-B, S2M-L and S2M-H as shown in Fig. 6(b). When employing SAM with ViT-B as the backbone in S2M, the best IoU is improved by 10.38% and 7.13% on the Static and L&F datasets than RPL, respectively. Comparatively, the IoU improvement from S2M-H to S2M-B is modest, with increases of 1.90% and 0.45% on Static and L&F. This pattern suggests that S2M remains robust across various SAM choices, owing to its reliance on box prompts for mask generation.
@@
AuIoU 25.10 35.32
Anomaly IoU 44.47 64.30
mean F1 36.35 41.53
AuIoU 4.12 47.04
Obstacle IoU 27.25 49.08
mean F1 6.27 48.38
AuIoU 15.51 28.07
RoadAnomaly IoU 33.80 32.96
mean F1 23.87 31.16
@@ 4.4 Efficiency Analysis
In our commitment to ensuring the model’s practical applicability in real-world scenarios, we also conducted an evaluation of its operational speed. Initially, we measured the standalone execution speed of RPL, followed by assessments of S2M-B, S2M-L, and S2M-H in Fig. 8. This evaluation involved processing all images in the Road Anomaly dataset and calculating the time taken for each image. The experiments were conducted on an NVIDIA RTX A5000 GPU. The runtime measurement encompassed the duration from the model receiving the input to producing the anomaly score or mask. For the RPL model, the processing time per image was recorded at 0.217s. Comparatively, S2M-B processed each image in 0.276s, showing only a marginal increase of 0.059s, which amounts to a 27.2% longer duration. The parameter of RPL model is 168M, and the total parameter of our S2M, built upon RPL, amounts to 300M. These results indicate that our method does not significantly burden resources during inference, thereby demonstrating its strong potential for practical deployment.
Figure 9: Failure cases for S2M. Challenges arise when OoD objects are too small for current detectors to effectively identify, leading to anomaly score maps lacking distinctive features. The left column shows original images, highlighting OoD objects with a magnifying glass, the right column showing the anomaly scores maps.

### v4

@@ Abstract
Semantic segmentation models, while effective for in-distribution categories, face challenges in real-world deployment due to encountering out-of-distribution (OoD) objects. Detecting these OoD objects is crucial for safety-critical applications. Existing methods rely on anomaly scores, but choosing a suitable threshold for generating masks presents difficulties and can lead to fragmentation and inaccuracy. This paper introduces a method to convert anomaly Score To segmentation Mask, called S2M, a simple and effective framework for OoD detection in semantic segmentation. Unlike assigning anomaly scores to pixels, S2M directly segments the entire OoD object. By transforming anomaly scores into prompts for a promptable segmentation model, S2M eliminates the need for threshold selection. Extensive experiments demonstrate that S2M outperforms the state-of-the-art by approximately 20% in IoU and 40% in mean F1 score, on average, across various benchmarks including Fishyscapes, Segment-Me-If-You-Can, and RoadAnomaly datasets. Code is available at https://github.com/WenjieZhao1/S2M.
@@ 1 Introduction
Semantic segmentation, a critical task in computer vision, bears substantial significance across various applications such as autonomous driving and aerial imagery analysis [18, 50]. While current semantic segmentation models demonstrate impressive performance, their practical deployment remains challenging. A significant obstacle lies in their limited ability to detect out-of-distribution (OoD) objects. Specifically, these models often assign pixels within the OoD object to one of the categories used in training, leading to inaccurate segmentation masks [32, 46, 26]. Addressing the OoD detection problem in semantic segmentation is important, given that inaccuracies in OoD objects masks can lead to erroneous conclusions, ultimately posing safety concerns in applications such as autonomous driving [50, 14, 38, 4, 3].
@@
Existing OoD detection methods in semantic segmentation address this issue by assigning an anomaly score for each pixel [24, 32, 46, 35, 2]. The pixels with high anomaly scores will be considered as part of OoD objects. The anomaly scores are typically derived by the probabilistic predictions of each pixel made by the segmentation model [22, 25, 36]. For example, PEBAL [46] proposes a pixel-wise energy-based segmentation method to compute the anomaly score. RPL [32] introduces a residual pattern learning module and computes the anomaly score by an energy-based method [46], effectively enhancing the model’s sensitivity to OoD pixels without compromising the segmentation performance on in-distribution (ID) data.
@@
While the anomaly score-based OoD detection methods can accurately identify anomalous pixels, they lack an effective way for segmenting the entire OoD object. In particular, to derive a mask for OoD objects, it is crucial to use a carefully chosen threshold to distinguish between anomalous and normal pixels [47, 9, 16]. However, determining the optimal threshold in practical applications can be a difficult task. In Fig. 2, it is evident that the optimal threshold range for existing OoD detection methods is quite narrow; even a slight deviation, either too high or too low, can result in inaccurate segmentation. Choosing the optimal threshold often requires a dedicated validation dataset for fine-tuning the threshold. In practice, such validation datasets are often not available. Even with the optimal threshold, the generated masks from anomaly score-based OoD detection methods may still require refinement, as the anomaly scores for certain pixels might be inaccurate, leading to fragmented or discontinuous masks. Fig. 1 illustrates various examples of masks generated by PEBAL [46] and RPL [32], two state-of-the-art anomaly score-based OoD detection methods. The fragmented masks can hardly be useful as the model users cannot accurately locate the OoD objects.
@@
In this paper, we present a method, named S2M, which converts anomaly Score To segmentation Mask, serving as a simple and general out-of-distribution (OoD) detection framework for semantic segmentation. Specifically, rather than deriving masks for OoD objects through thresholding anomaly scores, S2M aims to segment the entire OoD object. Thus S2M effectively mitigates the fragmentation issue associated with employing a threshold. In more details, given any anomaly scores generated by existing OoD detection methods, S2M generates box prompts from them using a prompt generator. The generated box prompt will approximately locate the OoD objects. Then, the box prompts are used as input for a promptable segmentation model to generate the masks for the OoD objects. Fig. 1 shows several examples of masks generated by S2M for the OoD objects. Different from using a threshold to generate the masks, the masks generated by S2M accurately segment the entire OoD objects while avoiding the inclusion of normal pixels in the mask. Compared with existing OoD detection methods, S2M has several advantages: 1) Simple: S2M is a simple pipeline that is easy to train and deploy, requiring no hyperparameter tuning. 2) General: S2M can be integrated with any anomaly score-based OoD detection methods to generate high-quality OoD masks. 3) Effective: S2M can accurately segment the OoD objects without creating fragmented and inaccurate masks.
@@
  • We extensively evaluate S2M on commonly used OoD segmentation benchmarks including Fishyscapes, Segment-Me-If-You-Can, and RoadAnomaly datasets, the results show that S2M improves the state-of-the-art by approximately 20% in IoU and 40% in mean F1 score, on average, across all benchmarks.
@@ 2 Related work
Recent trends in semantic segmentation have shifted towards transformer-based architectures and attention mechanisms, which offer improved handling of contextual relationships within images [13, 52, 45]. In contrast to these typical segmentation approaches, the prompt-based segmentation presents a unique paradigm. Its primary strength lies in its robust segmentation capabilities, enabling the efficient delineation of objects within an image. For example, CLIPSeg [34] introduced a system capable of generating image segmentation at test time based on any given prompt, whether text or image, thereby enabling a unified model for three distinct segmentation tasks. Segment Anything Model (SAM) [26] is known for its robust capability to generate precise masks for every object, demonstrating exceptional performance in object segmentation tasks.
@@ 3.2 Anomaly Score to Box Prompt
Figure 4: Box prompts can lead to more accurate segmentation for the OoD objects compared to point prompts. We derive point prompts from the locations corresponding to the extreme values of the anomaly scores. Visual analysis indicates that box prompts substantially improve the model’s tolerance to noise. We employ all generated box prompts to obtain a comprehensive mask of the OoD object, which ensures that the entire object is covered.
@@ 3.3 Outlier Exposure
In the OE process, a transformation function $T$ is applied to the OoD object masks $y^{out}$. It randomly rescales the OoD images and the corresponding masks at first. Then, it cuts or pads the images and masks into the dimension of the in-distribution images. The OE process is mathematically formulated as follows:
@@ 3.4 Prompt Generator Training
The initial step of training the prompt generator involves transforming the mask labels $y^{oe}$ into prompt labels. In particular, a transformation function $T_{box}$ is used to generate box prompt $B_{prompt}$ from $y^{oe}$ by finding the coordinates of the smallest rectangular region that can fully contain all the masked pixels.
@@ 3.5 Inference Pipeline
During inference, the input image is processed using a state-of-the-art OoD detection method, to compute an anomaly score map. This map is then fed into the prompt generator, which yields bounding box prompts indicating the locations of potential OoD objects. Subsequently, we employ a promptable segmentation model which takes both the prompts and the original image as inputs and produces masks of the OoD objects as shown in Fig. 5.
@@ 4 Experiments
DenseHybrid [17] 6.81 23.54 11.15 2.41 15.35 3.83 19.25 35.77 29.4 2.23 20.04 3.41 13.53 26.17 21.34
Synboost [11] 17.89 32.81 25.67 7.37 18.39 10.85 28.35 42.00 39.48 5.45 13.97 9.02 19.50 27.22 29.33
PEBAL [46] 8.21 26.92 13.31 1.49 6.35 2.57 23.98 42.39 35.09 0.57 6.74 1.06 15.51 33.80 23.87
RPL+CoroCL [32] 8.46 36.46 13.16 2.55 15.75 3.91 22.26 68.80 31.55 3.85 28.66 5.72 16.88 50.97 24.64
S2M (Ours) 64.78 69.99 70.24 29.30 30.46 35.31 54.33 77.54 60.40 58.12 67.64 64.96 54.31 58.49 61.66
Table 1: S2M outperforms the state-of-the-art methods on Fishyscapes Static (FS Static), Fishyscapes Lost and Found (FS Lost&Found), SMIYC Anomaly, SMIYC Obstacle and RoadAnomaly validation set. In our method, which is based on the anomaly score from RPL+CoroCL, the IoU is calculated without a threshold and by using all the produced prompt boxes. Both AuIoU and mean F1 scores are calculated by iterating through all thresholds. To facilitate a fair comparison with existing methods, we use the confidence of the box prompt as the anomaly score when calculating AuIoU and mean F1.
@@ 4.1 Experimental Setup
  • PEBAL [46]. This method introduces pixel-wise energy-biased abstention learning, which synergistically optimizes a novel pixel-wise anomaly abstention learning framework along with energy-based models.
@@ 4.2 Main Results
The results also show that existing anomaly score-based OoD detection methods, such as DenseHybrid, RPL, and PEBAL, are less effective in producing accurate masks for the OoD objects. While the RPL+CoroCL method achieves high IoU values, it does not perform as well across other metrics, such as the mean F1 score [5], revealing a shortfall in achieving highly accurate masks for OoD objects [5]. Our method, however, excels in all assessed metrics. The IoU achieved by our method surpasses that of the state-of-the-art methods by approximately 20% across all datasets, on average. Specifically, on the SMIYC-Obstacle validation dataset, our method outperforms RPL by a significant margin of 38.98%.
@@
0% 74.57 63.15 66.29 62.61 55.68 52.89
1% 77.54 60.40 67.64 64.96 58.49 61.66
2% 83.44 65.60 60.62 56.41 58.67 55.46
3% 74.89 59.92 56.53 48.11 56.18 54.93
4% 77.85 54.59 57.73 57.67 56.72 56.45
Table 2: Performance across different noise intensities. We introduced a random perturbation to the magnitude of the anomaly scores and observed that a fluctuation of 1% yielded the best training outcomes on these datasets. All experimental results reflect the performance after 80 epochs of training. The mean F1 scores are derived by systematically iterating through all thresholds in increments of 0.01.
@@ 4.3 Ablation Studies
Generalize to other anomaly scores. To further validate the generalization of our method, we leverage PEBAL [46] for generating the anomaly scores in S2M. Applying PEBAL to generate anomaly scores for the synthetic training dataset and utilizing the previously mentioned training strategy, we also observed similar performance improvement. Table 4 shows that a significant increase in terms of IoU, with improvements of 12.12% and 34.03% on the SMIYC anomaly and obstacle tracks [5] compared with PEBAL , respectively. This shows that S2M can generalize to various anomaly scores and be easily integrated with existing anomaly score-based OoD detection methods.
Data augmentation. We experimented with different levels of random fluctuations within a specific range to the values of the anomaly scores and subsequently tested the model’s performance on the Fishyscapes Static and Fishyscapes Lost & Found [3] datasets. Table 2 indicate that the model achieves optimal overall performance with the addition of 1% noise. The model exhibited a 2.97% increase in the best IoU on the SMIYC anomaly dataset, and there was also an improvement in the mean F1 scores across both datasets. On the SMIYC obstacle dataset, the model performance saw a slight improvement, with an increase of 1.35% in the best IoU. Furthermore, it was observed that an excessive noise level leads to a significant reduction in the mean F1 score, suggesting that overly strong noise can adversely affect model training, hindering the model’s ability to accurately generate prompts.
Different choices of SAM. We investigated various SAM backbone configurations, including ViT-B, ViT-L, and ViT-H [26], the corresponding S2M is denoted as S2M-B, S2M-L and S2M-H as shown in Fig. 7(b). When employing SAM with ViT-B as the backbone in S2M, the best IoU is improved by 33.53% and 14.71% on the Static and L&F datasets than RPL, respectively. Comparatively, the IoU improvement from S2M-H to S2M-B is modest, with increases of 0.89% and 0.03% on Static and L&F. This pattern suggests that S2M remains robust across various SAM choices, owing to its reliance on box prompts for mask generation.
@@
AuIoU 23.98 28.40
Anomaly IoU 42.39 54.51
mean F1 35.09 37.03
AuIoU 0.57 24.27
Obstacle IoU 6.74 40.77
mean F1 1.06 27.07
AuIoU 15.51 29.42
RoadAnomaly IoU 33.80 36.17
mean F1 23.87 33.30
@@ 4.4 Efficiency Analysis
For efficiency analysis, we first measured the standalone execution speed of RPL, followed by assessments of S2M-B, S2M-L, and S2M-H in Fig. 8. This evaluation involved processing all images in the Road Anomaly dataset and calculating the time taken for each image. The experiments were conducted on an NVIDIA RTX A5000 GPU. The runtime measurement encompassed the duration from the model receiving the input to producing the anomaly score or mask. For the RPL model, the processing time per image was recorded at 0.217s. Comparatively, S2M-B processed each image in 0.276s, showing only a marginal increase of 0.059s, which amounts to a 27.2% longer duration. The parameter of RPL model is 168M, and the total parameter of our S2M, built upon RPL, amounts to 300M. These results indicate that our method does not significantly burden resources during inference, thereby demonstrating its strong potential for practical deployment.
Figure 9: Failure cases for S2M. Challenges arise when OoD objects are too small for current detectors to identify, leading to anomaly score maps lacking distinctive features. The left column shows original images, highlighting OoD objects with a magnifying glass, the right column showing the anomaly scores maps.